OpenStudy (anonymous):

The graph of which quadratic equation has a vertex of (2, 1)? y = (x – 1)2 + 2 y = (x – 2)2 + 1 y = (x + 2)2 + 1 y = (x + 1)2 + 2

OpenStudy (campbell_st):

the vertex form of the equation is \[y = (x - h)^2 + k\] has a vertex at (h, k) so you need an equation when h = 2 and k = 1

OpenStudy (anonymous):

so it would be the third one? @campbell_st

OpenStudy (campbell_st):

nope... the 3rd one has h = -2 remember is (x - h) ( x - -2) = x + 2